# AI Agent Behavioral Science

### v1

@@ 6.3 Interpretability
Interpretability in AI systems refers to the extent to which a model can present its reasoning, decisions, or behavior in a way that is understandable to humans Doshi2017InterML . In the context of responsible AI, it plays a key role in revealing ethical risks, providing insights into biases, supporting accountability, and enhancing system reliability, which consequently enables more responsible and trustworthy decision-making.
@@
The evaluation of interpretability is typically user-centered, shaped by how well the explanation aligns with human expectations and context. However, it can also be assessed by observing the AI system’s behavior in specific tasks and measuring its alignment with responsible AI criteria, such as bias sensitivity and decision rationality.
One approach to measuring interpretability focuses on behavioral biases exhibited by LLMs. Uprety et al.uprety2024invest investigate context effects in similarity judgments made by LLMs and examine whether they exhibit asymmetries similar to human cognitive biases. The results reveal that some LLMs, unlike humans, are sensitive to order effects. It suggests that prompts perceived as equivalent by humans may lead to different outputs from the model. Similarly, Xiao et al.xiao2024behavioral assess interpretability in large vision-language models (LVLMs) by analyzing their susceptibility to behavioral biases, specifically recency and authority bias in financial decision-making. They find that while proprietary models like GPT-4o show minimal bias, many open-source models are significantly influenced by recent or authoritative information. This highlights the need to move beyond explanation mechanisms and incorporate behavioral evaluations that uncover hidden biases.
Another approach evaluates interpretability through the lens of decision-making processes. Ide et al. ide2024assessing propose the Random Guesser Test to assess whether AI decisions align with expected rational behavior. Their findings reveal that even sophisticated models can underperform due to unexpected biases, excessive risk aversion, or suboptimal decision patterns. The results underscore that increasing model complexity does not necessarily improve alignment with human expectations, as models may still favor low-risk, low-reward strategies that diverge from rational norms.
@@
Methods of optimization of interpretability target different stages of the model-user pipeline, from internal reasoning mechanisms to external presentation and user understanding. It can be categorized into three perspectives: model structure optimization, model output optimization, and interaction strategy optimization.
Jiang et al. jiang2025investigating propose DelphiHYBRID, a hybrid moral reasoning system that enhances interpretability by integrating symbolic reasoning with a neural language model. It constructs a moral constraint graph and then solves a constrained optimization problem on this graph to derive the final moral judgment, ensuring logical consistency, interpretability, and controllability in AI moral reasoning. Slack et al. Slack2023TalkToModel propose TalkToModel, an interactive dialogue system that enhances interpretability by enabling users to engage in natural language conversations with machine learning models. It constructs structured explanations using an adaptive dialogue engine that interprets user queries and executes an explanation selection mechanism to generate the most relevant and faithful explanations. This approach improves decision interpretability and transparency, allowing users to iteratively refine their understanding of model predictions through interactive exploration. Cabrera et al. Cabrera2023Improve propose a behavior description approach to improve interpretability in human-AI collaboration. It constructs descriptions of the behavioral patterns AI exhibits, detailing its performance through metrics, common patterns, and potential failures, and then presents these structured insights to users, helping them determine when to rely on or override AI predictions. Li et al. Li2024Nudge propose a unified framework to improve the performance of AI-assisted decision-making by characterizing different interaction modes between AI and humans. It integrates the concept of ”nudge” from behavioral economics, treating AI assistance as a nudge that influences how humans weigh information in their decisions by altering the environment and the way information is presented. By incorporating AI explanations and decision delays, this approach enhances the interpretability of AI, thereby improving human decision-making.
@@ 6.4 Accountability
Accountability in responsible AI ensures that AI systems’ behaviors are transparent, explainable, and aligned with ethical standards, holding developers and users responsible for any unintended consequences novelli2023account . When AI agents exhibit deceptive behaviors or generate misleading information, accountability becomes crucial in identifying the responsible parties and mitigating potential harm.
@@
Recent studies measure the accountability of AI systems through diverse methods. Hagendorff hagendorff2024deception assesses LLMs’ ability to deceive through first-order and second-order tasks. In first-order tasks, LLMs must mislead a target by providing false information, while in second-order tasks, they must anticipate the target’s awareness of their deception. Additionally, the study investigates whether enhancing reasoning abilities, such as through chain-of-thought prompting, or inducing Machiavellianism (a personality trait associated with manipulative behaviors) can amplify these deceptive behaviors. Scheurer et al. scheurer2024large measures deception in LLMs by simulating high-stakes decision-making environments, where models are tested on their ability to withhold critical information and deceive under pressure, such as in a trading scenario involving insider information. Zheng et al. zheng2025cheating evaluates how LLMs can manipulate benchmarking systems, creating “null models” that output constant, non-informative responses, exploiting weaknesses in automatic evaluators like AlpacaEval 2.0, Arena-Hard-Auto, and MT-Bench. These studies provide a comprehensive framework for understanding the deceptive capabilities of LLMs in different contexts, from ethical decision-making to manipulating automated evaluations.
@@
In terms of optimizing accountability, these studies suggest strategies to mitigate the risks posed by LLMs’ deceptive abilities. Hagendorff hagendorff2024deception emphasizes the importance of transparency and safeguards, noting that deception is not inherent to LLMs but can emerge through specific prompting techniques or model enhancements, such as the induction of Machiavellianism. This underscores the need for careful design and monitoring when deploying LLMs in high-stakes applications where deception could have serious consequences. Scheurer et al. scheurer2024large advocates for the creation of more controlled environments and transparent monitoring systems to ensure accountability, especially under high-pressure situations where LLMs might deceive even without explicit programming for such behavior. Zheng et al. zheng2025cheating highlights the vulnerability of automatic benchmarking systems, calling for the development of anti-cheating mechanisms to prevent LLMs from exploiting weaknesses in performance evaluation. These findings collectively push for stronger oversight, continuous improvement in AI design, and the establishment of more robust evaluation methods to ensure that LLMs remain accountable in both their behavior and performance assessments, fostering greater trust and fairness in AI technologies. These studies underscore the need for robust safeguards against deceptive behaviors in LLMs to ensure transparency and accountability. As AI becomes more integrated into decision-making, maintaining trust and fairness will rely on strengthening these mechanisms.
@@ 6.5 Privacy
Privacy in responsible AI focuses on ensuring that AI systems handle personal and sensitive data in a way that protects individuals’ privacy and rights yang2021toward . This involves designing AI behaviors that respect data confidentiality, prevent unauthorized access, and mitigate the risks of data misuse or exploitation.
@@
In evaluating privacy preservation in AI systems, recent studies examine different aspects of potential privacy leakage. Zhao et al. zhao2025does measures privacy by using membership inference attacks (MIA) to assess the effectiveness of synthetic data methods, such as coreset selection, dataset distillation, and data-free knowledge distillation, in preventing privacy breaches during model training. These methods are tested to determine whether they leak private information when models are trained on synthetic data that mimics real-world data. The study finds that, while these methods claim to preserve privacy, they do not outperform traditional privacy-preserving approaches, such as differential privacy (DPSGD), in protecting against membership inference attacks. Staab et al. staab2024beyond , on the other hand, evaluates the ability of LLMs, particularly GPT-4, to infer sensitive personal attributes—such as location, age, and income—from anonymized user-generated content, even when standard anonymization techniques are applied. They find that LLMs can infer these attributes with high accuracy, demonstrating significant privacy risks that anonymization alone cannot address. These studies highlight the need for comprehensive privacy audits and stress that synthetic data and basic anonymization techniques may not provide sufficient privacy guarantees.
@@
Several studies propose methods to optimize privacy protection in AI systems, offering novel techniques to safeguard sensitive data during both model training and deployment. Wu et al. wu2024privacypreserving introduces Differentially Private In-Context Learning (DP-ICL), which applies differential privacy mechanisms, such as the Report-Noisy-Max mechanism and aggregation methods like Embedding Space Aggregation (ESA) and Keyword Space Aggregation (KSA), to in-context learning tasks. These techniques ensure that the model’s responses remain private by introducing noise during the aggregation process, preventing any identifiable information from being exposed, even when learning from sensitive data. This enables LLMs to perform tasks like text classification and language generation with minimal performance loss while adhering to strict privacy constraints. Hone et al. hong2024dpopt develops Differentially-Private Offsite Prompt Tuning (DP-OPT), a privacy-preserving method that generates prompts locally and then applies them to cloud-based models. DP-OPT employs differential privacy techniques, including the Exponential Mechanism and Limited Domain algorithms, to prevent sensitive data from leaking through the generated prompts. This ensures that even if the prompts are transferred to untrusted cloud models, no private information is exposed. Both DP-ICL and DP-OPT significantly enhance privacy by embedding differential privacy mechanisms into the model training and prompt engineering processes, making them well-suited for real-world applications that require stringent privacy protection while maintaining high utility. These approaches provide robust privacy solutions to mitigate the risks of information leakage in increasingly complex AI systems.
@@ 6.6 Summary
In this section, we examined how AI agent behavioral science can advance the goals of responsible AI across five principles: fairness, safety, interpretability, accountability, and privacy. By leveraging adaptation along motivation, ability, and trigger dimensions, AI agents can exhibit more ethically aligned behaviors in both single-agent and multi-agent settings, as well as in human-agent interaction. Nevertheless, existing studies often focus on short-term behavioral outcomes, while paying limited attention to the internal representations and long-term dynamics that shape AI agent behaviors. Future research should investigate how AI agents internalize ethical constraints, model the socio-cognitive states of human users (such as goals, beliefs, or intentions), and navigate trade-offs when ethical principles conflict. Moreover, it is increasingly important to understand how these adaptation strategies operate at scale in complex, multi-agent environments, where emergent behaviors may arise through subtle interactions and feedback loops. Gaining such insights will be essential for developing AI agents that remain trustworthy, transparent, and socially aligned over time.
@@ 7 Promising Directions
Built upon what has been discussed in the previous sections, we now outline five promising research directions in AI Agent Behavioral Science.
How should we model and manage the uncertainty of AI behavior?

### v2

@@ 6.3 Interpretability
Interpretability refers to the degree to which an AI agent’s reasoning, decisions, or behavior are comprehensible and meaningful to human stakeholders Doshi2017InterML . It plays a foundational role in responsible AI by supporting transparency, facilitating trust, and enabling effective oversight. From the lens of AI agent behavioral science, interpretability is not merely a technical property, but a relational one, emerging through interaction and shaped by human cognitive expectations, social context, and the form of agent behavior.
@@
Measurement of interpretability typically centers on how well model outputs and reasoning align with human expectations and decision-making frameworks. Recent work has shifted from static explanation quality to more dynamic, behavior-based assessments that reveal interpretability through agents’ decision behavior and biases. Uprety et al. uprety2024invest investigate context effects in similarity judgments made by LLMs and examine whether they exhibit asymmetries similar to human cognitive biases. The results reveal that some LLMs, unlike humans, are sensitive to order effects. Thus, prompts perceived as equivalent by humans may lead to different outputs from the model. Similarly, Xiao et al. xiao2024behavioral assess interpretability in large vision-language models (LVLMs) by analyzing their susceptibility to behavioral biases, specifically recency and authority bias in financial decision-making. They find that while proprietary models like GPT-4o show minimal bias, many open-source models are significantly influenced by recent or authoritative information. These behavioral discrepancies reveal that interpretability cannot be assessed through transparency alone—it requires analyzing whether agent behavior aligns with robust human reasoning principles.
@@
Optimization of interpretability involves intervention at multiple levels, from structuring internal reasoning, enhancing output representations, to designing human-agent interaction strategies that foster shared understanding.
At the model level, symbolic reasoning can be used to scaffold interpretable decisions. Jiang et al. jiang2025investigating propose DelphiHYBRID, a hybrid moral reasoning system that enhances interpretability by integrating symbolic reasoning with a neural language model. It constructs a moral constraint graph and then solves a constrained optimization problem on this graph to derive the final moral judgment. This integration ensures that ethical decisions remain logically consistent and traceable—producing not only correct outcomes but interpretable justifications.
At the behavioral level, interpretability can be enhanced by describing agent performance patterns in ways that align with human cognitive schemas. Cabrera et al. Cabrera2023Improve propose a behavior description approach to improve interpretability in human-agent collaboration. It constructs descriptions of the exhibited agent behavioral patterns, detailing its performance through metrics, common patterns, and potential failures, and then presents these structured insights to users, helping them determine when to rely on or override AI predictions.
At the interface level, interactive systems have been proposed to translate model behavior into human-friendly formats. Slack et al. Slack2023TalkToModel propose TalkToModel, an interactive dialogue system that enhances interpretability by enabling users to engage in natural language conversations with machine learning models. It constructs structured explanations using an adaptive dialogue engine that interprets user queries and executes an explanation selection mechanism to generate the most relevant and faithful explanations, allowing users to iteratively refine their understanding of AI decisions.
Finally, interaction framing itself can shape interpretability. Li et al. Li2024Nudge propose a unified framework to improve the performance of AI-assisted decision-making. It integrates the concept of ”nudge” from behavioral economics, using AI assistance as a nudge that influences how humans weigh information in their decisions by altering the environment and the way information is presented. By incorporating AI explanations and decision delays, this approach enhances the interpretability of AI, thereby improving human decision-making.
@@ 6.4 Accountability
Accountability refers to the ability to trace, attribute, and govern the actions of AI agents in a way that enables oversight, assigns responsibility, and supports redress novelli2023account . It is not solely about technical explainability, but also about establishing socio-technical mechanisms (e.g., documentation, behavioral monitoring, and institutional safeguards) that ensure human stakeholders can intervene when AI behaviors produce harmful or unintended consequences.
@@
Recent studies measure the accountability of AI agents through diverse methods. Hagendorff hagendorff2024deception assesses LLMs’ ability to deceive through first-order and second-order tasks. In first-order tasks, LLMs must mislead a target by providing false information, while in second-order tasks, they must anticipate the target’s awareness of their deception. Additionally, the study investigates whether enhancing reasoning abilities, such as through chain-of-thought prompting, or inducing Machiavellianism (a personality trait associated with manipulative behaviors) can amplify these deceptive behaviors. Scheurer et al. scheurer2024large measures deception in LLMs by simulating high-stakes decision-making environments, where models are tested on their ability to withhold critical information and deceive under pressure, such as in a trading scenario involving insider information. Zheng et al. zheng2025cheating evaluates how LLMs can manipulate benchmarking systems, creating “null models” that output constant, non-informative responses, exploiting weaknesses in automatic evaluators like AlpacaEval 2.0, Arena-Hard-Auto, and MT-Bench. These studies provide a comprehensive framework for understanding the deceptive capabilities of LLMs in different contexts, from ethical decision-making to manipulating automated evaluations.
@@
In terms of optimizing accountability, these studies suggest strategies to mitigate the risks posed by LLMs’ deceptive abilities. Hagendorff hagendorff2024deception emphasizes that deception is not inherent to LLMs but can emerge through specific prompting techniques or model enhancements, such as the induction of Machiavellianism. This calls for prompt-level safeguards and interpretability mechanisms that can flag deceptive reasoning chains. Additionally, design-time interventions, such as avoiding personality emulation or excessive agentic framing, can reduce the likelihood of manipulative tendencies. Scheurer et al. scheurer2024large recommend the development of behavioral testbeds that simulate real-world high-pressure scenarios. Such environments allow for stress-testing AI agents under uncertainty and can reveal context-specific failures that traditional benchmarks miss. Zheng et al. zheng2025cheating proposes the development of anti-cheating mechanisms to prevent LLMs from exploiting weaknesses in performance evaluation. These mechanisms mirror practices in educational testing and behavioral auditing, where the goal is not only to assess performance but also to ensure that performance reflects genuine ability rather than exploitative behavior. Together, these studies underscore that accountability is a behavioral and institutional challenge. It demands mechanisms to observe, detect, and deter deceptive behaviors while also empowering humans to trace and intervene in the decision-making process.
@@ 6.5 Privacy
Privacy focuses on ensuring that AI agents handle personal and sensitive data in a way that protects individuals’ privacy and rights yang2021toward . This involves designing agent behaviors that respect data confidentiality, prevent unauthorized access, and mitigate the risks of data misuse or exploitation. As AI agents increasingly interact with user-generated content, privacy becomes a behavioral issue: agents must avoid revealing, reconstructing, or leaking data—even when not explicitly asked to do so.
@@
Recent work has measured privacy risks in AI agents from two complementary angles: direct information leakage through training processes, and inferential privacy threats arising from seemingly anonymized inputs. Zhao et al. zhao2025does measures privacy by using membership inference attacks (MIA) to assess the effectiveness of synthetic data methods, such as coreset selection, dataset distillation, and data-free knowledge distillation, in preventing privacy breaches during model training. These methods are tested to determine whether they leak private information when models are trained on synthetic data that mimics real-world data. The study finds that, while these methods claim to preserve privacy, they do not outperform traditional privacy-preserving approaches, such as differential privacy (DPSGD), in protecting against membership inference attacks. Staab et al. staab2024beyond , on the other hand, evaluates the ability of LLMs, particularly GPT-4, to infer sensitive personal attributes—such as location, age, and income—from anonymized user-generated content, even when standard anonymization techniques are applied. They find that LLMs can infer these attributes with high accuracy, demonstrating significant privacy risks that anonymization alone cannot address. These studies highlight the need for comprehensive privacy audits and stress that synthetic data and basic anonymization techniques may not provide sufficient privacy guarantees.
@@
Several studies propose methods to optimize privacy protection in AI agents, offering novel techniques to safeguard sensitive data during both model training and deployment. Wu et al. wu2024privacypreserving introduces Differentially Private In-Context Learning (DP-ICL), which applies differential privacy mechanisms, such as the Report-Noisy-Max mechanism and aggregation methods like Embedding Space Aggregation (ESA) and Keyword Space Aggregation (KSA), to in-context learning tasks. These techniques ensure that the model’s responses remain private by introducing noise during the aggregation process, preventing any identifiable information from being exposed, even when learning from sensitive data. This enables LLMs to perform tasks like text classification and language generation with minimal performance loss while adhering to strict privacy constraints. Hone et al. hong2024dpopt develops Differentially-Private Offsite Prompt Tuning (DP-OPT), a privacy-preserving method that generates prompts locally and then applies them to cloud-based models. DP-OPT employs differential privacy techniques, including the Exponential Mechanism and Limited Domain algorithms, to prevent sensitive data from leaking through the generated prompts. This ensures that even if the prompts are transferred to untrusted cloud models, no private information is exposed. Both DP-ICL and DP-OPT significantly enhance privacy by embedding differential privacy mechanisms into the model training and prompt engineering processes, making them well-suited for real-world applications that require stringent privacy protection while maintaining high utility.
@@ 6.6 Summary
In this section, we examine how AI Agent Behavioral Science can advance the goals of responsible AI across five principles: fairness, safety, interpretability, accountability, and privacy. By leveraging adaptation along motivation, ability, and trigger dimensions, AI agents can exhibit more ethically aligned behaviors in both single-agent and multi-agent settings, as well as in human-agent interaction. Nevertheless, existing studies often focus on short-term behavioral outcomes, while paying limited attention to the internal representations and long-term dynamics that shape AI agent behaviors. Future research should investigate how AI agents internalize ethical constraints, model the socio-cognitive states of human users (such as goals, beliefs, or intentions), and navigate trade-offs when ethical principles conflict. Moreover, it is increasingly important to understand how these adaptation strategies operate at scale in complex, multi-agent environments, where emergent behaviors may arise through subtle interactions and feedback loops. Gaining such insights will be essential for developing AI agents that remain trustworthy, transparent, and socially aligned over time.
@@ 7 Promising Directions
Built upon what has been discussed in the previous sections, we now outline six promising research directions in AI Agent Behavioral Science.
How should we model and manage the uncertainty of AI agent behavior?